Dan bought 5 1/4 gallons of gas on Friday and ​ 6 1/2 ​ gallons of gas on Saturday. How many liters of gas did he buy? Assume 1 gallon = 3.8 liters.

Enter your answer, as a decimal, in the box.
1
Expert's answer
2019-11-21T10:44:50-0500

"5\frac{1}{4}+6\frac{1}{2}=\frac{21}{4}+\frac{26}{4}=\frac{47}{4}."

"\frac{47}{4} \cdot 3.8=44.65 \;l."